Suppose that $60,000 is invested at 6% interest, compounded annually. After time t, in years, it grows to the amount A given by the function A(t)=$60,000(1.06)^t. Find the amount of time after which there will be $300,000 in the account.
Answer by ewatrrr(24785)

 
Hi
A(t)=$60,000(1.06)^t
Find the amount of time after which there will be $300,000 in the account
300000/60000 = (1.06)^t
5 = (1.06)^t
log5/log(1.06) = t
27.62 yrs = t